Given the following demand function and total cost function as:

Q = 250-5P and TC = 0.5Q2 + 5Q + 125 

Calculate the equilibrium price, Equilibrium Quantity and the total profit

A.        under perfectly competitive market structure.

B.         Under monopolistic competitive market. 


1
Expert's answer
2022-01-05T15:21:45-0500

A. P=MC

"MC=TC' =( 0.5Q2 + 5Q + 125 )'=Q+5"

"P=\frac{250-Q}{5}"

"\frac{250-Q}{5}=Q+5"

250-Q=5Q+25

225=6Q

Q=37.5

"P=\frac{250-37.5}{5}=42.5"

"Profit=TR-TC=PQ-(0.5Q2 + 5Q + 125)=37.5\times42.5-(0.5(37.5)^2+5\times37.5+125)=578.125"


B.

MR=MC

MR-MC=0

MC=Q+5

"TR=\frac{250-Q}{5}\times Q=\frac{250Q-Q^2}{5}="

"MR=TR'=(\frac{250Q-Q^2}{5})'=50-\frac{2Q}{5}"

"50-\frac{2Q}{5}=Q+5"

"45=Q+\frac{2Q}{5}"

225=5Q+2Q

225=7Q

Q=32.14


"P=\frac{250-Q}{5}=\frac{250-32.14}{5}=43.572"

"Profit=TR-TC=PQ-(0.5Q2 + 5Q + 125)=32.14\times43.572-(0.5(32.14)^2+5\times32.14+125)=598.2143"